# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v2

@@ II-E Summary and Lessons Learned
  • Beyond general-purpose systems like Windows and Linux, OS for embodied LM agents should cater to robotic environments such as UAVs, autonomous vehicles, and robots. Integrating systems such as ROS (Robot Operating System) for machine-centric applications will be key to ensure seamless functionality and promote embodied intelligence.
@@
TABLE IV: A Summary of LLM Stages, Utilized Technologies, and Representives
@@ III Large Model Agent Networks: Working Principles
In this section, we first explore the general architecture of LM agent networks, covering their definition and underlying engine. We then discuss the key characteristics and communication modes of LM agent networks. Next, we introduce the cloud-edge-end cooperative architecture for networking LM agents, followed by an in-depth analysis of interaction strategies, including cooperation, partial cooperation, and competition. Finally, we examine cooperation paradigms in LM agent networks, focusing on data, computation, and knowledge sharing.
@@ III-A2 Engine of LM Agent Networks
  • Foundation models, such as LLMs, LVMs, and large multimodal models (LMMs), serve as the brains of LM agents that comprehensively power them in planning, action, memory, interaction, and security capacities, as detailed in Sect. II-C. Table IV summarizes the basic stages of LLM services. (i) Advanced reasoning: LM technologies empower AI agents with advanced reasoning abilities such as multi-persona self-planning and grounded planning. For instance, CoT [49] and ToT [50] reasoning allow LM agents to break down complex tasks into manageable sub-tasks. (ii) Few/zero-shot generalization: Pretrained on extensive corpora, LMs demonstrate few-shot and zero-shot generalization capabilities [61], allowing LM agents to transfer knowledge seamlessly between tasks. (iii) Tool usage capability: Through the use of various tools, LM agents can gather and integrate valuable information from diverse sources to execute and adapt actions under complex scenarios. (iv) Adaptability: Through continuous learning and adaptation, LM agents can accumulate knowledge over time by learning from new data and experiences [77].
@@
TABLE V: A Summary of Intra-Agent and Inter-Agent Communications for Connected LM Agents
@@ III-B Key Characteristics of LM Agent Networks
3) Semantic-aware Communication. Unlike traditional data-centric communication paradigms, LM agents leverage semantic-aware communication, prioritizing the exchange of meaningful or contextually relevant information for reduced communication overheads. Semantic-aware intra-agent or inter-agent communications facilitate a richer contextual understanding of the environment with improved accuracy of predictions and actions.
  • Intra-agent Semantic-aware Communication. Intra-agent communication focuses on synchronizing the brain and body of LM agents. These communications is semantic-aware, meaning only contextually relevant data such as changes in the environment are transmitted to minimize bandwidth usage. For instance, an LM-powered household robot sends its brain in the cloud only the changes in the environment and its position, rather than transmitting the entire environmental data repeatedly.
@@ III-C Communication Mode of LM Agents
Every LM agent consists of two parts: (i) the LM-empowered cyber brain located in the cloud, edge servers, or end devices and (ii) the corresponding physical or software-form body. Every LM agent can actively interact with other LM agents, the virtual/real environment, and humans. For connected LM agents, there exist two typical communication modes: intra-agent communications for seamless data/knowledge synchronization between brain and physical body within an LM agent, and inter-agent communications for efficient coordination between LM agents. Table V summarizes the comparison of the two communication modes.
@@
For instance, traditional static APIs are highly efficient and portable but lack versatility, while natural language communication offers high versatility and portability but often sacrifices efficiency. LM agents bridge this gap by understanding, manipulating, and responding to other agents through natural language while integrating external tools, writing code, and invoking APIs. LM agents excel at following instructions, including implementing routines through code. Furthermore, they autonomously negotiate protocols and reach consensus on strategies and behaviors in complex scenarios. Marro et al. propose the Agora protocol [25] to addresses the communication trilemma by employing standardized routines for frequent communications, natural language for infrequent communications (e.g., negotiation and error handling), and LLM-written routines for scenarios that fall between them.
In addition to Agora, other agent communication protocols have been developed, e.g., Agent Protocol (adopted by AutoGPT), SmythOS, and Azure AI Agent Service. Selecting or designing an appropriate agent communication protocol requires careful consideration of task scenarios (e.g., wired backhaul or radio access networks) and task-specific requirements (e.g., robustness, latency, and cost constraints). Additionally, adapting these protocols to diverse systems and agent platforms presents significant challenges, highlighting the need for flexible and interoperable solutions.
@@ III-D2 Cloud-Edge-End Cooperative Framework Design
LM lightweighting techniques, e.g., knowledge distillation [26], quantization [27], pruning [28], and hardware acceleration [29], can be employed to deploy compact LM variants tailored to edge/end computational capacities. For example, VPGTrans [26] addresses the high computational costs of training visual prompt generators for multimodal LLMs by significantly reducing GPU hours and training data needed via transferring an existing visual prompt generator from one multimodal LLM to another. LLM-Pruner [28] compresses LLMs without compromising their multi-task solving and language generation abilities, using a task-agnostic structural pruning method that significantly reduces the need for extensive retraining data and computational resources. Hardware accelerators, such as GPUs and TPUs, can enhance computational efficiency and provide robust support for cloud training and edge deployment. Agile-Quant [29] enhances the efficiency of LLMs on edge devices through activation-guided quantization and hardware acceleration by utilizing a SIMD-based 4-bit multiplier and efficient TRIP matrix multiplication, achieving up to 2.55x speedup while maintaining high task performance.
@@ III-D3 Advantages
By assigning specialized roles to cloud, edge, and end devices, the cloud-edge-end architecture enhances flexibility to accommodate varying task complexity and resource availability. For instance, computation-intensive LM tasks can be offloaded to cloud servers, less demanding LM tasks can be handled by edge servers to reduce latency and bandwidth usage, while end devices can offload latency-sensitive LM operations to edge servers. Additionally, hierarchical collaborative learning is essential for LM agents to acquire global knowledge across large geographical areas, with LM updates of edge/end agents synchronized centrally in the cloud. This hierarchical collaboration across cloud, edge, and end devices offers several key advantages:
@@
  • Security and Privacy: Localized processing at the end layer reduces the exposure of sensitive data, while the cloud and edge nodes provides robust security mechanisms for cross-layer interactions.
@@ III-E Interaction Strategies of Connected LM Agents
The multi-agent interactions can be categorized into distributed cooperation, distributed competition, and partial cooperation [82, 63], each of which involves different strategies to jointly optimize the collective or individual outcomes.
@@ III-E1 Distributed Cooperation
a) Distributed Cooperative LM Fine-tuning for LM Agents. LMs, typically designed for general tasks, require fine-tuning with vast domain-specific data on edges to provide agent services in specialized fields. However, due to their massive size, fine-tuning LMs demands substantial computational power and energy, as shown in Table IV. To fine-tune LMs on heterogeneous, resource-constrained edge environments while ensuring privacy protection, we introduce two paradigms: collaborative split learning (SL) and federated edge learning, as below.
@@
  • Cooperative SL. SL distributes partitioned LMs on edge devices and servers. In SL, as shown in Fig. 13(a), edge devices upload data features to edge servers for co-inference, reducing local computation while retaining major processing tasks at the edge. By placing only a submodel on edge devices, SL minimizes their computational burden and can integrate with parameter-efficient fine-tuning (PEFT) techniques [91] to further reduce workload. However, deploying SL for LM agents in edge-end environments introduces several challenges. First, communication costs for transmitting high-dimensional data can be significant. For example, partitioning large models generates smashed data that should be uploaded, resulting in substantial overhead. In the case of GPT-3 Medium and 100 samples (1024 tokens each), this can reach approximately 400 MB per training round [92]. Second, privacy concerns arise due to target token leakage. In common edge SL configurations, the input module is placed on the edge and the output module on the server, requiring edge devices to upload target tokens for training [93]. This can expose sensitive user data and compromise privacy. U-shaped SL [94] retains the initial and final neural layers or transformer blocks on edge devices, placing only intermediate layers on edge servers, thereby safeguarding label privacy to a large extent.
@@
  • Federated edge learning. In cloud-edge collaborative LM training, LMs are initially pre-trained in the cloud, utilizing its extensive computational resources and large-scale datasets. Federated fine-tuning technology is then performed on edge devices using domain-specific data from various edge servers, enabling LM agent services tailored to specific requirements of diverse localized downstream tasks and optimizing resource utilization across the network. However, due to the massive size of LMs (often containing billions of parameters), uploading model updates from numerous edge servers for aggregation during the federated fine-tuning process imposes significant communication overheads and operational costs. Additionally, periodic LM local fine-tuning at the edge consumes substantial computational resources, which often exceed the capacity of resource-constrained edge nodes. The offsite-tuning transfer learning technique, in conjunction with PEFT techniques [91], such as Prefix-tuning and LoRA, enable edge devices to update and transmit only a subset of model parameters, thereby significantly reducing both communication and computational overheads [85]. Specifically, as shown in Fig. 14, the cloud employs offsite-tuning transfer learning technique to generate a lightweight proxy model that encodes task-specific knowledge with a minimal number of parameters. Subsequently, edge nodes fine-tune the PEFT module of this proxy model to further reduce the communication and computational costs.
@@
b) Distributed Cooperative LM Inference for LM Agents. On-server LM inference offloads computation to cloud servers but compromises user privacy by requiring raw data uploads. On-device inference, on the other hand, preserves privacy but imposes heavy computational burden on edge devices. To resolve this dilemma in resource-constrained environments, mixture-of-experts (MoE) and split inference are two key approaches. Additionally, LM agents can leverage role-playing for cooperative inference in general applications.
  • Cooperative LM inference via MoE. MoE enhances large models (LMs) by partitioning them into specialized expert networks, each tailored to process specific input types [31]. A gating network dynamically assigns tokens to the most appropriate expert(s), enabling sparse activation that reduces unnecessary computations and improves efficiency. MoE has been shown to boost performance by 45% while consuming only one-third of the resources required by dense models [95]. In cloud-edge architectures, as shown in Fig. 15, cloud servers store the expert models, and edge servers and devices download only the relevant experts for local inference. This method optimizes resource utilization, accelerates processing, and makes the deployment of large models viable in resource-constrained environments. For instance, EdgeMoE [96] stores non-expert weights in memory while placing larger expert weights on disk, activating only the necessary experts. This approach reduces memory consumption by 2.6 to 3.2 times compared to full model loading. However, a series of challenges remain in MoE edge deployment, as follows. 1) MoE models are primarily designed for high-performance cloud data centers equipped with powerful GPUs and ample network bandwidth. However, when deployed on edge devices, MoE models must accommodate the wide variety of hardware capabilities, ranging from basic IoT devices to advanced edge servers. This heterogeneity requires flexibility in MoE design to ensure high performance while adapting to the varying resource constraints of edge environments. 2) Edge environments are highly dynamic, which complicates inference tasks. When local expert networks are unavailable, they must be dynamically downloaded, a process influenced by network speed and device capacity. Additionally, selecting the appropriate expert network adds complexity to model verification and management. 3) As the expert networks required for inference tasks may also expose users’ environmental context and behavioral patterns, it is essential to incorporate privacy protection measures when designing the inference engine.
@@
  • Cooperative split inference. Split inference, as outlined in the 3GPP 5G specifications [89], partitions an LM between edge devices and edge/cloud servers, offloading major part of computation to edge/cloud servers for mitigated device-side workload. It offers a middle ground between on-server and on-device inference. As shown in Fig. 13(b), edge devices perform initial inference using the user-side submodel and upload intermediate features to the server for further processing, thereby optimizing resource usage while preserving privacy by keeping sensitive raw data on the device. According to [46], split inference for LLMs can be implemented in two ways: 1) In encoder-decoder LMs such as BART [97], the encoder is placed on the device, while the more computationally demanding decoder is on the server. 2) In decoder-only LMs such as GPT-series, lightweight components (e.g., embeddings, early Transformer blocks) stay on the device, while heavier components are offloaded to the server. Additionally, critical feed-forward network (FFN) parameters can be stored on the device to save memory. However, split inference faces challenges such as high communication costs due to large data uploads, latency from reliance on both edge devices and servers, and privacy concerns from potential leakage of intermediate outputs. Early exit strategies help lower computational latency on edge devices and servers. Transformer outputs can further be compressed via quantization, pruning, and merging [27]. Additionally, progressive split inference [98] can eliminate unnecessary transmission of intermediate token representations while maintaining desired inference accuracy.
  • Role-playing. Role-playing [99, 67] is a typical method for facilitating cooperation among LM agents. 1) Instruction-following cooperation. Li et al. [100] propose a cooperative agent framework that employs role-playing with inception prompting to guide agents autonomously toward task completion. The system begins with human-provided ideas and role assignments, refined by a task-specifier agent. An AI user and an AI assistant then collaborate through multi-turn dialogues, with the AI user instructing and the assistant responding until the task is completed. 2) Instruction-free cooperation. Fig. 16 shows an example of role-playing based novel writing using collaborative agents [101]. ChatDev [67] is a virtual software company operated by “software agents” with diverse roles, e.g., chief officers, programmers, and test engineers. Following the waterfall model, it divides development into design, coding, testing, and documentation, breaking each phase into atomic subtasks managed by a chat chain. To create cooperative agents with adaptive behaviors, ProAgent [84] assigns each agent three roles: planner, verificator, and controller. The planner formulates high-level skills, incorporating belief revision and knowledge management. The verificator assesses the feasibility of skills, identifies issues if a skill fails, and initiates re-planning when needed. If the skill is feasible, the controller decomposes it into executable actions. Park et al. [99] create a community of 25 generative agents in a sandbox world named Smallville, where agents are represented by sprite avatars. These agents perform daily tasks, form opinions, and interact, mimicking human-like behaviors.
@@
Figure 16: An example of role-playing for novel writing using collaborative agents [101]. In the draft stage, the Planner creates an initial agent team and action plan, refining it through feedback from the Agent and Plan Observers. The Action Observer assigns tasks, monitors execution, and adjusts the plan based on performance. Agents collaboratively refine and execute tasks, adapting as needed.
@@
  • Edge LM Caching. Due to the limited storage capacity, edge servers need to optimize the cache hit ratio by placing popular LMs on edge servers. For instance, TrimCaching [102] optimizes LLM model placement by leveraging LLM parameter sharing to maximize cache efficiency under storage and latency constraints. By storing a single copy of shared parameter blocks on each edge server, it enhances storage utilization. Practical edge LM caching deployment also needs to consider user mobility patterns and inter-edge collaboration and competition strategies under different LM partitioning methods (e.g., MoE and SL).
  • Edge LM Replacement. As user demands evolve, previously cached LMs may no longer match current request patterns. In such cases, edge servers should replace older models with more relevant ones. This process incurs high communication costs, as transferring large-scale models can impose substantial overhead on mobile backhaul networks. Recency-based and frequency-based are two classic strategies, replacing the least recently used (LRU) or least frequently used (LFU) items with updated ones. Popularity-based methods, such as SwapMoE [103], prioritize caching highly activated expert networks, with importance determined by activation frequency during inference tasks. However, these approaches often neglect shared parameter blocks across LLMs and the potential for cooperative caching among edge nodes. Distributed reinforcement learning can be leveraged to make dynamic replacement decisions without requiring complete information from other edge nodes.
  • Edge LM Routing. To further minimize service latency, it is essential to optimize edge LM routing strategies that direct users to the most suitable edge server based on network conditions, queue length, and model availability. Additionally, due to the mobility of agents, it requires seamless model migration across edge servers to maintain optimal service delivery. Further research efforts are required for intelligently managing user requests and distributing LM workloads across edge nodes to enhance performance, minimize latency, and maximize resource utilization.
  • Edge Caching for RAG. LM agents’ performance is often tightly linked to the quality and relevance of the knowledge they possess. Therefore, selecting which data to cache for LM updates via RAG is essential to ensure that LM agents provide reliable, up-to-date services. Efficient knowledge cache management not only improves response times but also supports the continuous learning and adaptation of LM agents in dynamic environments. Further research is needed to tackle the joint LM caching and knowledge caching problem within the context of RAG under practical constraints of edge environments.

### v3

@@ II-E Summary and Lessons Learned
  • Beyond general-purpose OSs (e.g., Windows and Linux), OS for embodied LM agents should cater to robotic environments such as UAVs, autonomous vehicles, and robots. Integrating systems such as ROS (Robot Operating System) for machine-centric communications will be key to ensure seamless functionality and promote embodied intelligence.
@@
  • The capabilities of LM agents have expanded significantly, enabling them to perform a wider range of tasks with enhanced efficiency. However, most existing virtual agent prototypes operate within isolated ecosystems and simulate agents only on a single device. Besides, existing virtual/embodied agent prototypes struggle with cost-effectiveness, cross-domain adaptability, and scalability, hindering their broader adoption. Addressing these challenges requires further research and innovation to enhance real-world implementation.
Figure 14: The general architecture of LM agent networks in bridging human, physical, and cyber worlds. LM agent has five constructing modules: planning, action, memory, interaction, and security modules. The engine of connected LM agents is empowered by a combination of five cutting-edge technologies: foundation models, knowledge-related technologies, interaction, digital twin, and multi-agent collaboration.
TABLE V: A Summary of LLM Stages, Utilized Technologies, and Representives
@@ III Large Model Agent Networks: Working Principles
In this section, we first explore the general architecture of LM agent networks, covering their definition and underlying engine. We then discuss the key characteristics and communication modes of LM agent networks. Next, we introduce a cloud-edge-end cooperative framework for networking LM agents, followed by an in-depth analysis of interaction strategies, including cooperation, partial cooperation, and competition. Finally, we examine cooperation paradigms in LM agent networks, focusing on data, computation, and knowledge sharing. Fig. 13 shows the organization structure of this section.
@@ III-A General Architecture of LM Agent Networks
As shown in Fig. 14, LM agent networks bridge the human, physical, and cyber worlds. Specifically, LM agents interact with humans through HMI technologies such as NLP, with the assistance of handheld devices and wearables, to understand human intentions, desires, and beliefs. LM agents can synchronize data and statuses between the physical body and the digital brain through digital twin technologies, as well as perceive and act upon the surrounding virtual/real environment. LM agents can be interconnected in the cyber space through efficient cloud-edge networking for efficient sharing of data, knowledge, and computation results to promote multi-agent collaboration.
@@ III-A2 Engine of LM Agent Networks
  • Foundation models, such as LLMs, large vision models (LVMs), and VLMs, serve as the brains of LM agents that comprehensively power them in planning, action, memory, interaction, and security capacities, as detailed in Sect. II-C. Table V summarizes the basic stages of LLM services. (i) Advanced reasoning: LM technologies empower AI agents with advanced reasoning abilities such as multi-persona self-planning and grounded planning. For instance, CoT [44] and ToT [45] reasoning allow LM agents to break down complex tasks into manageable sub-tasks. (ii) Few/zero-shot generalization: Pretrained on extensive corpora, LMs demonstrate few-shot and zero-shot generalization capabilities [57], allowing LM agents to transfer knowledge seamlessly between tasks. (iii) Tool usage capability: Through the use of various tools, LM agents can gather and integrate valuable information from diverse sources to execute and adapt actions under complex scenarios. (iv) Adaptability: Through continuous learning and adaptation, LM agents can accumulate knowledge over time by learning from new data and experiences [79].
@@
TABLE VI: A Summary of Intra-Agent and Inter-Agent Communications for Connected LM Agents
@@ III-B Key Characteristics of LM Agent Networks
3) Semantic-aware Communication. Unlike traditional communication on the Internet prioritizing reliable data transmission, agent-based communication is fundamentally task-oriented, which exhibits three unique features [89]:
  • Computing-oriented communications: Agent interactions prioritize transmitting information that directly supports computational workflows of tasks. Agent communications achieve this by first interpreting data semantics (e.g., via knowledge base alignment) and then transmitting only task-critical abstractions.
  • Persistent communications: Agent collaborations often span extended durations, requiring continuous adaptation to evolving task states. Agent communications leverage persistent memory to maintain shared context (e.g., incremental knowledge updates), enabling agents to reference prior interactions efficiently.
  • Memory-based communications: Agent tasks frequently exhibit sequential dependencies (e.g., iterative problem-solving). Agent communications exploit memory to compress information hierarchically. For instance, transmitting only the changes relative to previously shared knowledge.
Semantic-aware intra-agent and inter-agent communications facilitate a richer contextual understanding of the environment with improved accuracy of predictions and actions.
  • Intra-agent semantic-aware communication. Intra-agent communication focuses on brain-body synchronization of LM agents. These communications are semantic-aware, meaning only contextually relevant data are transmitted to minimize bandwidth usage. For instance, an LM-powered household robot sends its brain in the cloud only the changes in the environment and its position, rather than transmitting the entire environmental data repeatedly.
@@ III-C Communication Mode of LM Agents
Every LM agent consists of two parts: (i) the LM-empowered cyber brain located in the cloud, edge servers, or end devices and (ii) the corresponding physical or software-form body. Every LM agent can actively interact with other LM agents, the virtual/real environment, and humans. For connected LM agents, there exist two typical communication modes: intra-agent communications for seamless data/knowledge synchronization between brain and physical body within an LM agent, and inter-agent communications for efficient coordination between LM agents. Table VI summarizes the comparison of the two communication modes.
@@
For instance, traditional static APIs are highly efficient and portable but lack versatility, while natural language communication offers high versatility and portability but often sacrifices efficiency. LM agents bridge this gap by understanding, manipulating, and responding to other agents through natural language while integrating external tools, writing code, and invoking APIs. LM agents excel at following instructions, including implementing routines through code. Furthermore, they autonomously negotiate protocols and reach consensus on strategies and behaviors in complex scenarios. Marro et al. propose the Agora protocol [19] to address the communication trilemma by employing standardized routines for frequent communications, natural language for infrequent communications (e.g., negotiation and error handling), and LLM-written routines for scenarios that fall between them.
In addition to Agora, other agent communication protocols have been developed, e.g., Model Context Protocol (MCP) [92] and Agent Network Protocol (ANP) [93]. Selecting or designing an appropriate agent communication protocol requires careful consideration of task scenarios (e.g., wired backhaul or radio access networks) and task-specific requirements (e.g., robustness, latency, and cost constraints). Additionally, adapting these protocols to diverse systems and agent platforms presents significant challenges, highlighting the need for flexible and interoperable solutions.
@@ III-D1 Illustrating Example in Autonomous Driving
Figure 16: The taxonomy of interaction strategies of LM agent networks.
@@ III-D2 Cloud-Edge-End Cooperative Framework Design
LM lightweighting techniques, e.g., knowledge distillation [20], quantization [21], pruning [22], and hardware acceleration [23], can be employed to deploy compact LM variants tailored to edge/end computing capacities. For instance, VPGTrans [20] addresses the high computational costs of training visual prompt generators for multimodal LLMs by significantly reducing GPU hours and training data needed via transferring an existing visual prompt generator from one multimodal LLM to another. LLM-Pruner [22] compresses LLMs without compromising their multi-task solving and language generation abilities, using a task-agnostic structural pruning method that significantly reduces the need for extensive retraining data and computational resources. Hardware accelerators, e.g., GPUs and TPUs, enhance computational efficiency and provide robust support for cloud training and edge deployment. Agile-Quant [23] enhances the efficiency of LLMs on edge devices via activation-guided quantization and hardware acceleration by utilizing a SIMD-based 4-bit multiplier and efficient TRIP matrix multiplication, achieving up to 2.55x speedup while maintaining high task performance.
@@ III-D3 Advantages
By assigning specialized roles to cloud, edge, and end devices, the cloud-edge-end architecture enhances flexibility to accommodate varying task complexity and resource availability. For instance, compute-intensive LM tasks can be offloaded to cloud servers, less demanding LM tasks can be handled by edge servers to reduce latency and bandwidth usage, while end devices can offload latency-sensitive LM operations to edge servers. Additionally, hierarchical collaborative learning is essential for LM agents to acquire global knowledge across large geographical areas, with LM updates of edge/end agents synchronized centrally in the cloud. This hierarchical collaboration across cloud, edge, and end devices offers several key advantages:
@@
  • Security and Privacy: Localized processing at the end layer reduces the exposure of sensitive data, while the cloud and edge nodes provide robust security mechanisms for cross-layer interactions.
@@ III-E Interaction Strategies of Connected LM Agents
The multi-agent interactions can be categorized into distributed cooperation, distributed competition, and partial cooperation [84, 59], each of which involves different strategies to jointly optimize the collective or individual outcomes. Fig. 16 illustrates the taxonomy of interaction strategies of LM agent networks.
@@ III-E1 Distributed Cooperation
a) Distributed Cooperative LM Fine-tuning for LM Agents. LMs, typically designed for general tasks, require fine-tuning with vast domain-specific data on edges to provide agent services in specialized fields. However, due to their massive size, fine-tuning LMs demands substantial computational power and energy, as shown in Table V. To fine-tune LMs on heterogeneous, resource-constrained edge environments while ensuring privacy protection, we introduce two paradigms: collaborative split learning (SL) and federated edge learning, as below.
@@
① Cooperative SL. SL distributes partitioned LMs on edge devices and servers. In SL, as shown in Fig. 17(a), edge devices upload data features to edge servers for co-inference, reducing local computation while retaining major processing tasks at the edge. By placing only a submodel on edge devices, SL minimizes their computational burden and can integrate with parameter-efficient fine-tuning (PEFT) techniques [96] to further reduce workload. However, deploying SL for LM agents in edge-end environments introduces several challenges.
  • Communication costs for transmitting high-dimensional data can be significant. For instance, partitioning LMs generates smashed data that should be uploaded, resulting in substantial overhead. In the case of GPT-3 Medium and 100 samples (1024 tokens each), this can reach approximately 400 MB per training round [97].
  • Privacy concerns arise due to target token leakage. In common edge SL configurations, the input module is placed on the edge and the output module on the server, requiring edge devices to upload target tokens for training [98]. This can expose sensitive user data and compromise privacy. U-shaped SL [99] retains the initial and final neural layers or transformer blocks on edge devices, placing only intermediate layers on edge servers, thereby safeguarding label privacy to a large extent.
@@
② Federated edge learning. In cloud-edge collaborative LM training, LMs are initially pre-trained in the cloud, leveraging extensive computational resources and large-scale datasets. Federated fine-tuning technology is then performed on edge devices using domain-specific data from various edge servers, enabling LM agent services tailored to localized downstream tasks while optimizing resource utilization across the network. However, due to the enormous size of LMs (often containing billions of parameters), transmitting model updates from numerous edge servers for aggregation during federated fine-tuning imposes significant communication overheads and operational costs. Additionally, periodic local fine-tuning at the edge consumes substantial computational resources, often exceeding the capacity of resource-constrained edges.
The offsite-tuning transfer learning technique, in conjunction with PEFT techniques [96], such as Prefix-tuning and LoRA, enables edge devices to update and transmit only a subset of model parameters, significantly reducing both communication and computational overheads [87]. As shown in Fig. 18, the cloud generates a lightweight proxy model using offsite-tuning transfer learning, encoding task-specific knowledge with a minimal number of parameters. Edge nodes then fine-tune the PEFT module of this proxy model to further reduce communication and computational costs.
@@
b) Distributed Cooperative LM Inference for LM Agents. On-server LM inference offloads computation to cloud servers but compromises user privacy by requiring raw data uploads. On-device inference, on the other hand, preserves privacy but imposes a heavy computational burden on edge devices. To resolve this dilemma in resource-constrained environments, mixture-of-experts (MoE) and split inference are two key approaches. Additionally, LM agents can leverage role-playing for cooperative inference in general applications.
① Cooperative LM inference via MoE. MoE enhances LMs by partitioning them into specialized expert networks, each tailored to process specific input types [25]. A gating network dynamically assigns tokens to the most appropriate expert(s), enabling sparse activation that reduces unnecessary computations and improves efficiency. MoE has been shown to boost performance by 45% while consuming only one-third of the resources required by dense models [100]. In cloud-edge architectures, as shown in Fig. 19, cloud servers store the expert models, and edge servers and devices download only the relevant experts for local inference. This method optimizes resource utilization, accelerates processing, and makes the deployment of LMs viable under resource-constrained environments. For instance, EdgeMoE [101] stores non-expert weights in memory while placing larger expert weights on disk, activating only the necessary experts. This approach reduces memory consumption by 2.6 to 3.2 times compared to full model loading. However, a series of challenges remain in edge deployment, as follows.
  • Heterogeneous edge devices: MoE models are primarily designed for high-performance cloud data centers equipped with powerful GPUs and ample network bandwidth. However, when deployed on edge devices, MoE models must accommodate the wide variety of hardware capabilities, ranging from basic IoT devices to advanced edge servers. This heterogeneity requires flexibility in MoE design to ensure high performance while adapting to the varying resource constraints of edge environments.
  • Dynamic edge inference: Edge environments are highly dynamic, which complicates inference tasks. When local expert networks are unavailable, they must be dynamically downloaded, a process influenced by network speed and device capacity. Additionally, selecting the appropriate expert network adds complexity to model verification and management.
  • Privacy disclosure: As the expert networks required for inference tasks may also expose users’ environmental context and behavioral patterns, it is essential to incorporate privacy protection measures when designing the inference engine.
@@
② Cooperative split inference. Split inference, as outlined in the 3GPP 5G specifications [94], partitions an LM between edge devices and edge/cloud servers, offloading major part of computation to edge/cloud servers for mitigated device-side workload. It offers a middle ground between on-server and on-device inference. As shown in Fig. 17(b), edge devices perform initial inference using the user-side submodel and upload intermediate features to the server for further processing, thereby optimizing resource usage while preserving privacy by keeping sensitive raw data on the device. According to [39], split inference for LLMs can be implemented in two ways:
  • In encoder-decoder LMs such as BART [102], the encoder is placed on the device, while the more computationally demanding decoder is on the server.
  • In decoder-only LMs such as GPT-series, lightweight components (e.g., embeddings, early Transformer blocks) stay on the device, while heavier components are offloaded to the server. Additionally, critical feed-forward network (FFN) parameters can be stored on the device to save memory.
However, split inference faces challenges such as high communication costs due to large data uploads, latency from reliance on both edge devices and servers, and privacy concerns from potential leakage of intermediate outputs. Early exit strategies help lower computational latency on edge devices and servers. Transformer outputs can further be compressed via quantization, pruning, and merging [21]. Additionally, progressive split inference [103] can eliminate unnecessary transmission of intermediate token representations while maintaining desired inference accuracy.
@@
Figure 20: An example of role-playing for novel writing with collaborative agents [104]. In the draft stage, the Planner assembles an initial agent team and action plan, refining it via feedback from Agent and Plan Observers. The Action Observer assigns tasks, monitors execution, and adjusts the plan based on performance. Agents collaboratively execute tasks, adapting as needed.
③ Role-playing. Role-playing [105, 63] is a typical method for facilitating cooperation among LM agents.
  • Instruction-following cooperation. Li et al. [106] present a cooperative agent architecture that employs role-playing combined with inception prompting to guide LLM agents autonomously toward effective task completion. The system begins with human-provided ideas and role assignments, refined by a task-specifier agent. An AI user and an AI assistant then collaborate through multi-turn dialogues, with the AI user instructing and the assistant responding until the task is completed.
  • Instruction-free cooperation. Fig. 20 shows an example of role-playing based novel writing using collaborative agents [104]. ChatDev [63] is a virtual software company operated by “software agents” with diverse roles, e.g., chief officers, programmers, and test engineers. Following the waterfall model, it divides development into design, coding, testing, and documentation, breaking each phase into atomic subtasks managed by a chat chain. To create cooperative agents with adaptive behaviors, ProAgent [86] assigns each agent three roles: planner, verificator, and controller. The planner formulates high-level skills, incorporating belief revision and knowledge management. The verificator assesses the feasibility of skills, identifies issues if a skill fails, and initiates re-planning when needed. If the skill is feasible, the controller decomposes it into executable actions. Park et al. [105] create a community of 25 generative agents in a sandbox world named Smallville, where agents are represented by sprite avatars. These agents perform daily tasks, form opinions, and interact, mimicking human-like behaviors.
@@
① Edge LM caching. Due to the limited storage capacity, edge servers need to optimize the cache hit ratio by placing popular LMs on edge servers. For instance, TrimCaching [107] optimizes LLM model placement by leveraging LLM parameter sharing to maximize cache efficiency under storage and latency constraints. By storing a single copy of shared parameter blocks on each edge server, it enhances storage utilization. Practical edge LM caching deployment also needs to consider user mobility patterns and inter-edge collaboration and competition strategies under different LM partitioning methods (e.g., MoE and SL).
② Edge LM replacement. As user demands evolve, previously cached LMs may no longer match current request patterns. In such cases, edge servers should replace older models with more relevant ones. This process incurs high communication costs, as transferring large-scale models can impose substantial overhead on mobile backhaul networks. Recency-based and frequency-based are two classic strategies, replacing the least recently used (LRU) or least frequently used (LFU) items with updated ones. Popularity-based methods, such as SwapMoE [108], prioritize caching highly activated expert networks, with importance determined by activation frequency during inference tasks. However, these approaches often neglect shared parameter blocks across LLMs and the potential for cooperative caching among edge nodes. Distributed reinforcement learning can be leveraged to make dynamic replacement decisions without requiring complete information from other edge nodes.
③ Edge LM routing. To further minimize service latency, it is essential to optimize edge LM routing strategies that direct users to the most suitable edge server based on network conditions, queue length, and model availability. Additionally, due to the mobility of agents, it requires seamless model migration across edge servers to maintain optimal service delivery. Further research efforts are required for intelligently managing user requests and distributing LM workloads across edge nodes to enhance performance, minimize latency, and maximize resource utilization.
④ Edge caching for RAG. LM agents’ performance is often tightly linked to the quality and relevance of the knowledge they possess. Therefore, selecting which data to cache for LM updates via RAG is essential to ensure that LM agents provide reliable, up-to-date services. Efficient knowledge cache management not only improves response times but also supports the continuous learning and adaptation of LM agents in dynamic environments. Further research is needed to tackle the joint LM caching and knowledge caching problem within the context of RAG under practical constraints of edge environments.